ON THE QUANTITATIVE ANALYSIS OF RESONANCE RAMAN SPECTRA: A SYSTEMATIC INVESTI-GATION OF HYDROGEN BONDING IN ELECTRONICALLY-EXCITED ACETYLACETONE

\underline{S. A. BROADBENT}, C. CHATTERJEE, and P. H. VACCARO, Department of Chemistry, Yale University, P.O. Box 208107, New Haven, CT 06520; B. R. JOHNSON, Department of Chemistry, Rice University, P.O. Box 1892, Houston, TX 77251.

The cis-enol tautomers of simple β-diketones such as acetylacetone (H_{3}C-CO-CH_{2}-CO-CH_{3}) are ideal target compounds for the investigation of hydrogen bonding and proton transfer, exhibiting a variety of intramolecular processes (e.g., low-barrier hydrogen bonding) that have been predicted to play pivotal roles in the behavior of substantially larger complexes. Resonance Raman (RR) spectroscopy has been used to probe the electronically-excited B̃^{1}B_{2} (π^{∗}π) potential energy surface of acetylacetone, thereby elucidating the changes in structure and dynamics that accompany π^{∗} ← π electron promotion of the isolated (vapor-phase) species. Data acquired at discrete excitation wavelengths spanning the B̃−X̃ absorption system (λ_{max} ≈ 262 nm) displayed pronounced differences in intensity patterns. The selective activity of overtone and combination bands involving displacement of the H–chelated ring indicated a low-barrier, hydrogen-bonding motif for the B̃^{1}B_{2} manifold. The comprehensive interpretation of all experimental findings was facilitated by ab initio geometry optimizations and force-field calculations performed for the pertinent electronic states at substantial levels of coupled-cluster theory (CCSD and EOM–CCSD with augmented correlation-consistent basis sets). The Hessian matrix and gradient vector computed for the electroncially-excited surface at the fully-relaxed ground-state (X̃^{1}A_{1}) geometry led to a harmonically-extrapolated B̃^{1}B_{2} equilibrium structure that bears evidence for the low-barrier hydrogen bonding phenomenon. The vibrational results emerging from this “vertical Hessian” treatment were employed as initial parameters for a least-squares regression procedure designed to simulate observed RR spectra by means of a time-dependent propagator formalism that incorporated effects arising from the Duschinsky rotation of normal coordinates and the non-Condon character of transition moments. Quantitative information extracted from these analyses will be discussed, with particular emphasis directed towards unraveling the unimolecular dynamics of acetylacetone.